# Performance Analysis and Power Allocation for Massive MIMO ISAC Systems

### v2

@@ B-B ZF Beamforming
$=\left|\mathbb{E}\left\{\left(\hat{{\mathbf{h}}}_{k}+{\mathbf{e}}_{k}\right)^{\scriptscriptstyle\mathsf{H}}\left(\sqrt{\gamma_{k}}\check{{\mathbf{h}}}_{k}+\sqrt{\eta_{k}}{\mathbf{a}}\right)\right\}\right|^{2}=\gamma_{k}\left|\mathbb{E}\left\{\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}\breve{{\mathbf{h}}}_{k}\right\}\right|^{2}$
@@
$E_{0}$ $=\mathbb{E}\left\{\left|\sqrt{\gamma_{k}}\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}\breve{{\mathbf{h}}}_{k}+\sqrt{\eta_{k}}\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}{\mathbf{a}}+{\mathbf{e}}_{k}^{\scriptscriptstyle\mathsf{H}}(\sqrt{\gamma_{k}}\breve{{\mathbf{h}}}_{k}+\sqrt{\eta_{k}}{\mathbf{a}})\right|^{2}\right\}$
$=\gamma_{k}\mathbb{E}\left\{|\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}\breve{{\mathbf{h}}}_{k}|^{2}\right\}+\eta_{k}\mathbb{E}\left\{|\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}{\mathbf{a}}|^{2}\right\}$
$\quad+\sqrt{\gamma_{k}\eta_{k}}\mathbb{E}\left\{\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}\breve{{\mathbf{h}}}_{k}\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}{\mathbf{a}}\right\}$
$\quad+\eta_{k}\mathbb{E}\left\{\left|{\mathbf{e}}_{k}^{\scriptscriptstyle\mathsf{H}}{\mathbf{a}}\right|^{2}\right\}+\gamma_{k}\mathbb{E}\left\{\left|{\mathbf{e}}_{k}^{\scriptscriptstyle\mathsf{H}}\breve{{\mathbf{h}}}_{k}\right|^{2}\right\},$ (73)
@@
$\mathbb{E}\left\{|\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}{\mathbf{a}}|^{2}\right\}=\mathbb{E}\left\{\lVert\hat{{\mathbf{h}}}_{k}\rVert^{2}\right\}=N_{\mathtt{t}}\xi_{k},$ (75)
$\mathbb{E}\left\{\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}\breve{{\mathbf{h}}}_{k}\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}{\mathbf{a}}\right\}=\mathbb{E}\left\{\Big{[}\hat{{\mathbf{H}}}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{H}}}(\hat{{\mathbf{H}}}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{H}}})^{-1}\Big{]}_{kk}\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}{\mathbf{a}}\right\}=\bm{0},$ (76)
$\mathbb{E}\left\{\left|{\mathbf{e}}_{k}^{\scriptscriptstyle\mathsf{H}}{\mathbf{a}}\right|^{2}\right\}=\mathbb{E}\left\{\left\lVert{\mathbf{e}}_{k}\right\rVert^{2}\right\}=N_{\mathtt{t}}\epsilon_{k}.$ (77)
Here, (75)–(77) follow from the fact that all the elements of ${\mathbf{a}}$ are deterministic with unit modulus and $\hat{{\mathbf{h}}}_{k}\sim\mathcal{CN}(0,\xi_{k}\textbf{{I}}_{N_{\mathtt{t}}})$. To compute $\mathbb{E}\{|{\mathbf{e}}_{k}^{\scriptscriptstyle\mathsf{H}}\breve{{\mathbf{h}}}_{k}|^{2}\}$, we write $\breve{{\mathbf{h}}}_{k}=\hat{{\mathbf{H}}}(\hat{{\mathbf{H}}}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{H}}})^{-1}{\mathbf{i}}_{k}$, where ${\mathbf{i}}_{k}$ is the $k$th column of $\textbf{{I}}_{K}$. As a result, we have
$\mathbb{E}\left\{|{\mathbf{e}}_{k}^{\scriptscriptstyle\mathsf{H}}\breve{{\mathbf{h}}}_{k}|^{2}\right\}=\epsilon_{k}\mathbb{E}\left\{\lVert\breve{{\mathbf{h}}}_{k}\rVert^{2}\right\}=\epsilon_{k}\mathbb{E}\left\{\text{trace}\left(\breve{{\mathbf{h}}}_{k}\breve{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}\right)\right\}$
$=\epsilon_{k}\mathbb{E}\left\{\text{trace}\left((\hat{{\mathbf{H}}}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{H}}})^{-1}{\mathbf{i}}_{k}{\mathbf{i}}_{k}^{\scriptscriptstyle\mathsf{H}}(\hat{{\mathbf{H}}}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{H}}})^{-1}\hat{{\mathbf{H}}}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{H}}}\right)\right\}$
@@
$\mathbb{E}\left\{\left|\mathtt{UI}_{kj}\right|^{2}\right\}=\gamma_{j}\mathbb{E}\left\{|\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}\breve{{\mathbf{h}}}_{j}|^{2}\right\}+\eta_{j}\mathbb{E}\left\{|\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}{\mathbf{a}}|^{2}\right\}$
$\qquad+\eta_{j}\mathbb{E}\left\{|{\mathbf{e}}_{k}^{\scriptscriptstyle\mathsf{H}}{\mathbf{a}}|^{2}\right\}+\gamma_{j}\mathbb{E}\left\{|{\mathbf{e}}_{k}^{\scriptscriptstyle\mathsf{H}}\breve{{\mathbf{h}}}_{j}|^{2}\right\}.$
@@ Appendix C Proof of Theorem 2
C-A CRLB for MRT
From (15) and (36), we have
${\mathbf{R}}_{\mathtt{MRT}}$ $=\mathbb{E}\left\{\hat{{\mathbf{H}}}\bm{\Gamma}^{2}\hat{{\mathbf{H}}}^{\scriptscriptstyle\mathsf{H}}+\rho{\mathbf{a}}{\mathbf{a}}^{\scriptscriptstyle\mathsf{H}}+{\mathbf{a}}\bar{\bm{\eta}}^{\scriptscriptstyle\mathsf{T}}\bm{\Gamma}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{H}}}^{\scriptscriptstyle\mathsf{H}}+\hat{{\mathbf{H}}}\bm{\Gamma}\bar{\bm{\eta}}{\mathbf{a}}^{\scriptscriptstyle\mathsf{H}}\right\}$
$=\sum_{k=1}^{K}\gamma_{k}\mathbb{E}\left\{\hat{{\mathbf{h}}}_{k}\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}\right\}+\rho{\mathbf{a}}{\mathbf{a}}^{\scriptscriptstyle\mathsf{H}}=\bm{\xi}^{\scriptscriptstyle\mathsf{T}}\bm{\gamma}\textbf{{I}}_{N_{\mathtt{t}}}+\rho{\mathbf{a}}{\mathbf{a}}^{\scriptscriptstyle\mathsf{H}}.$ (90)
From (29), (35), and (88)–(90), and noting that $\left\lVert{\mathbf{a}}\right\rVert^{2}=N_{\mathtt{t}}$ and $\left\lVert{\mathbf{b}}\right\rVert^{2}=N_{\mathtt{r}}$, we have
$J_{\theta\theta}$ $=\kappa\left|\alpha\right|^{2}\left(\bm{\xi}^{\scriptscriptstyle\mathsf{T}}\bm{\gamma}\left(N_{\mathtt{r}}\left\lVert\dot{{\mathbf{a}}}_{\theta}\right\rVert^{2}+N_{\mathtt{t}}\lVert\dot{{\mathbf{b}}}_{\theta}\rVert^{2}\right)+\rho N_{\mathtt{t}}^{2}\lVert\dot{{\mathbf{b}}}_{\theta}\rVert^{2}\right),$ (91)
$J_{\phi\phi}\!$ $=\!\kappa\!\left|\alpha\right|^{2}\left(\bm{\xi}^{\scriptscriptstyle\mathsf{T}}\bm{\gamma}\left(N_{\mathtt{r}}\left\lVert\dot{{\mathbf{a}}}_{\phi}\right\rVert^{2}\!+\!N_{\mathtt{t}}\lVert\dot{{\mathbf{b}}}_{\phi}\rVert^{2}\right)\!+\rho N_{\mathtt{t}}^{2}\lVert\dot{{\mathbf{b}}}_{\phi}\rVert^{2}\!\right)\!,$ (92)
$J_{\theta\phi}\!$ $=\!\kappa\left|\alpha\right|^{2}\!\!\left(\!\bm{\xi}^{\scriptscriptstyle\mathsf{T}}\bm{\gamma}\!\!\left(N_{\mathtt{r}}\dot{{\mathbf{a}}}_{\theta}^{\scriptscriptstyle\mathsf{H}}\dot{{\mathbf{a}}}_{\phi}+N_{\mathtt{t}}\dot{{\mathbf{b}}}_{\theta}^{\scriptscriptstyle\mathsf{H}}\dot{{\mathbf{b}}}_{\phi}\right)\!+\!\rho N_{\mathtt{t}}^{2}\dot{{\mathbf{b}}}_{\theta}^{\scriptscriptstyle\mathsf{H}}\dot{{\mathbf{b}}}_{\phi}\right)\!,$ (93)
${\mathbf{J}}_{\bar{\bm{\alpha}}\bar{\bm{\alpha}}}$ $=\kappa\left(\bm{\xi}^{\scriptscriptstyle\mathsf{T}}\bm{\gamma}N_{\mathtt{t}}N_{\mathtt{r}}+\rho N_{\mathtt{t}}^{2}N_{\mathtt{r}}\right)\textbf{{I}}_{2}.$ (94)
Based on (35) and (87)–(89) we have $\dot{{\mathbf{G}}}_{\theta}^{\scriptscriptstyle\mathsf{H}}{\mathbf{G}}=({\mathbf{a}}\dot{{\mathbf{b}}}_{\theta}^{\scriptscriptstyle\mathsf{H}}+\dot{{\mathbf{a}}}_{\theta}{\mathbf{b}}^{\scriptscriptstyle\mathsf{H}}){\mathbf{b}}{\mathbf{a}}^{\scriptscriptstyle\mathsf{H}}=\bm{0}$ and $\dot{{\mathbf{G}}}_{\phi}^{\scriptscriptstyle\mathsf{H}}{\mathbf{G}}=({\mathbf{a}}\dot{{\mathbf{b}}}_{\phi}^{\scriptscriptstyle\mathsf{H}}+\dot{{\mathbf{a}}}_{\phi}{\mathbf{b}}^{\scriptscriptstyle\mathsf{H}}){\mathbf{b}}{\mathbf{a}}^{\scriptscriptstyle\mathsf{H}}=\bm{0}$, yielding
${\mathbf{J}}_{\theta\bar{\bm{\alpha}}}=\bm{0},\ {\mathbf{J}}_{\phi\bar{\bm{\alpha}}}=\bm{0},\ \tilde{J}_{\phi\bar{\bm{\alpha}}}=J_{\phi\phi}.$ (95)
Substituting (91)–(95) into (37) and (38), we obtain (39) and (40) with $\bm{\xi}_{\mathtt{bf}}=\bm{\xi}_{\mathtt{MRT}}$ as defined in Theorem 2 .
C-B CRLB with ZF
Similar to (90), for the ZF precoder, we have
${\mathbf{R}}_{\mathtt{ZF}}$ $=\underbrace{\mathbb{E}\left\{\hat{{\mathbf{H}}}(\hat{{\mathbf{H}}}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{H}}})^{-1}\bm{\Gamma}^{2}(\hat{{\mathbf{H}}}(\hat{{\mathbf{H}}}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{H}}})^{-1})^{\scriptscriptstyle\mathsf{H}}\right\}}_{\triangleq{\mathbf{A}}}+\rho{\mathbf{a}}{\mathbf{a}}^{\scriptscriptstyle\mathsf{H}}.$
@@ C-B CRLB with ZF
Let ${\mathbf{A}}={\mathbf{V}}\bm{\Lambda}{\mathbf{V}}^{\scriptscriptstyle\mathsf{H}}$ be the eigenvalue decomposition of ${\mathbf{A}}$. Then ${\mathbf{U}}{\mathbf{V}}\bm{\Lambda}{\mathbf{V}}^{\scriptscriptstyle\mathsf{H}}{\mathbf{U}}^{\scriptscriptstyle\mathsf{H}}={\mathbf{A}}$, implying that any set of mutually orthogonal vectors constitutes valid eigenvectors for ${\mathbf{A}}$. Thus, ${\mathbf{A}}$ must be a scaled identity, i.e., ${\mathbf{A}}=c\textbf{{I}}_{N_{\mathtt{t}}}$. As a result, we can write ${\mathbf{R}}_{x,\mathtt{ZF}}=c\textbf{{I}}_{N_{\mathtt{t}}}+\rho{\mathbf{a}}{\mathbf{a}}^{\scriptscriptstyle\mathsf{H}}$, where $c=\frac{1}{N_{\mathtt{t}}(N_{\mathtt{t}}-K)}\sum_{k=1}^{K}\frac{\gamma_{k}}{\xi_{k}}=\bm{\xi}_{\mathtt{ZF}}^{\scriptscriptstyle\mathsf{T}}\bm{\gamma}$, with $\bm{\xi}_{\mathtt{ZF}}$ defined in Theorem 2.
Now, by applying derivations similar to those for MRT, we obtain
$J_{\theta\theta}$ $=\kappa\left|\alpha\right|^{2}\left(\bm{\xi}_{\mathtt{ZF}}^{\scriptscriptstyle\mathsf{T}}\bm{\gamma}\left(\left\lVert\dot{{\mathbf{a}}}_{\theta}\right\rVert^{2}+\lVert\dot{{\mathbf{b}}}_{\theta}\rVert^{2}\right)+\rho\lVert\dot{{\mathbf{b}}}_{\theta}\rVert^{2}\right),$ (97)
@@
$J_{\theta\phi}$ $=\kappa\left|\alpha\right|^{2}\left(\bm{\xi}_{\mathtt{ZF}}^{\scriptscriptstyle\mathsf{T}}\bm{\gamma}\left(\dot{{\mathbf{a}}}_{\theta}^{\scriptscriptstyle\mathsf{H}}\dot{{\mathbf{a}}}_{\phi}+\dot{{\mathbf{b}}}_{\theta}^{\scriptscriptstyle\mathsf{H}}\dot{{\mathbf{b}}}_{\phi}\right)+\rho\dot{{\mathbf{b}}}_{\theta}^{\scriptscriptstyle\mathsf{H}}\dot{{\mathbf{b}}}_{\phi}\right),$ (98)
@@
$J_{\phi\phi}$ $=\kappa\left|\alpha\right|^{2}\left(\bm{\xi}_{\mathtt{ZF}}^{\scriptscriptstyle\mathsf{T}}\bm{\gamma}\left\lVert\dot{{\mathbf{a}}}_{\phi}\right\rVert^{2}+\lVert\dot{{\mathbf{b}}}_{\phi}\rVert^{2}\right)+\rho\lVert\dot{{\mathbf{b}}}_{\phi}\rVert^{2},$ (99)
@@
${\mathbf{J}}_{\bar{\bm{\alpha}}\bar{\bm{\alpha}}}$ $=\kappa\left(\bm{\xi}_{\mathtt{ZF}}^{\scriptscriptstyle\mathsf{T}}\bm{\gamma}+\rho\right)\textbf{{I}}_{2},$ (100)

### v3

@@ B-B ZF Beamforming
$=\left|\mathbb{E}\left\{\left(\hat{{\mathbf{h}}}_{k}+{\mathbf{e}}_{k}\right)^{\scriptscriptstyle\mathsf{H}}\left(\sqrt{\gamma_{k}}\check{{\mathbf{h}}}_{k}+\sqrt{\eta_{k}}{{\mathbf{v}}}\right)\right\}\right|^{2}=\gamma_{k}\left|\mathbb{E}\left\{\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}\breve{{\mathbf{h}}}_{k}\right\}\right|^{2}$
@@
$E_{0}$ $=\mathbb{E}\left\{\left|\sqrt{\gamma_{k}}\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}\breve{{\mathbf{h}}}_{k}+\sqrt{\eta_{k}}\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}{{\mathbf{v}}}+{\mathbf{e}}_{k}^{\scriptscriptstyle\mathsf{H}}(\sqrt{\gamma_{k}}\breve{{\mathbf{h}}}_{k}+\sqrt{\eta_{k}}{{\mathbf{v}}})\right|^{2}\right\}$
$=\gamma_{k}\mathbb{E}\left\{|\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}\breve{{\mathbf{h}}}_{k}|^{2}\right\}+\eta_{k}\mathbb{E}\left\{|\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}{{\mathbf{v}}}|^{2}\right\}$
$\quad+\sqrt{\gamma_{k}\eta_{k}}\mathbb{E}\left\{\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}\breve{{\mathbf{h}}}_{k}\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}{{\mathbf{v}}}\right\}$
$\quad+\eta_{k}\mathbb{E}\left\{\left|{\mathbf{e}}_{k}^{\scriptscriptstyle\mathsf{H}}{{\mathbf{v}}}\right|^{2}\right\}+\gamma_{k}\mathbb{E}\left\{\left|{\mathbf{e}}_{k}^{\scriptscriptstyle\mathsf{H}}\breve{{\mathbf{h}}}_{k}\right|^{2}\right\},$ (80)
@@
$\mathbb{E}\left\{|\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}{{\mathbf{v}}}|^{2}\right\}=\mathbb{E}\left\{\lVert\hat{{\mathbf{h}}}_{k}\rVert^{2}\right\}=N_{\mathtt{t}}\xi_{k},$ (82)
$\mathbb{E}\left\{\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}\breve{{\mathbf{h}}}_{k}\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}{{\mathbf{v}}}\right\}=\mathbb{E}\left\{\Big{[}\hat{{\mathbf{H}}}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{H}}}(\hat{{\mathbf{H}}}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{H}}})^{-1}\Big{]}_{kk}\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}{{\mathbf{v}}}\right\}=\bm{0},$ (83)
$\mathbb{E}\left\{\left|{\mathbf{e}}_{k}^{\scriptscriptstyle\mathsf{H}}{{\mathbf{v}}}\right|^{2}\right\}=\mathbb{E}\left\{\left\lVert{\mathbf{e}}_{k}\right\rVert^{2}\right\}=N_{\mathtt{t}}\epsilon_{k}.$ (84)
Here, (82)–(84) follow from the fact that all the elements of ${{\mathbf{v}}}$ are deterministic with unit modulus and $\hat{{\mathbf{h}}}_{k}\sim\mathcal{CN}(0,\xi_{k}\textbf{{I}}_{N_{\mathtt{t}}})$. To compute $\mathbb{E}\{|{\mathbf{e}}_{k}^{\scriptscriptstyle\mathsf{H}}\breve{{\mathbf{h}}}_{k}|^{2}\}$, we write $\breve{{\mathbf{h}}}_{k}=\hat{{\mathbf{H}}}(\hat{{\mathbf{H}}}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{H}}})^{-1}{\mathbf{i}}_{k}$, where ${\mathbf{i}}_{k}$ is the $k$th column of $\textbf{{I}}_{K}$. As a result, we have
$\mathbb{E}\left\{|{\mathbf{e}}_{k}^{\scriptscriptstyle\mathsf{H}}\breve{{\mathbf{h}}}_{k}|^{2}\right\}=\epsilon_{k}\mathbb{E}\left\{\lVert\breve{{\mathbf{h}}}_{k}\rVert^{2}\right\}=\epsilon_{k}\mathbb{E}\left\{\mathtt{trace}\left(\breve{{\mathbf{h}}}_{k}\breve{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}\right)\right\}$
$=\epsilon_{k}\mathbb{E}\left\{\mathtt{trace}\left((\hat{{\mathbf{H}}}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{H}}})^{-1}{\mathbf{i}}_{k}{\mathbf{i}}_{k}^{\scriptscriptstyle\mathsf{H}}(\hat{{\mathbf{H}}}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{H}}})^{-1}\hat{{\mathbf{H}}}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{H}}}\right)\right\}$
@@
$\mathbb{E}\left\{\left|\mathtt{UI}_{kj}\right|^{2}\right\}=\gamma_{j}\mathbb{E}\left\{|\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}\breve{{\mathbf{h}}}_{j}|^{2}\right\}+\eta_{j}\mathbb{E}\left\{|\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}{{\mathbf{v}}}|^{2}\right\}$
$\qquad+\eta_{j}\mathbb{E}\left\{|{\mathbf{e}}_{k}^{\scriptscriptstyle\mathsf{H}}{{\mathbf{v}}}|^{2}\right\}+\gamma_{j}\mathbb{E}\left\{|{\mathbf{e}}_{k}^{\scriptscriptstyle\mathsf{H}}\breve{{\mathbf{h}}}_{j}|^{2}\right\}.$
@@ Appendix C Proof of Theorem 2
The equalities in (39) and (40) follow directly from the fact that the CRLBs for $\theta$ and $\phi$ correspond to the first and second diagonal entries of ${\mathbf{J}}_{\theta,\phi}$ in (37), i.e., $\widetilde{\mathtt{CRLB}}_{\mathtt{bf},\theta}(\bm{\gamma},\rho)=[{\mathbf{J}}_{\theta,\phi}^{-1}]_{11}$ and $\widetilde{\mathtt{CRLB}}_{\mathtt{bf},\phi}(\bm{\gamma},\rho)=[{\mathbf{J}}_{\theta,\phi}^{-1}]_{22}$, where, $[{\mathbf{A}}]_{ij}$ denotes the $(i,j)$-th entry of matrix ${\mathbf{A}}$. In the following, we prove the closed form expressions for $J_{\theta\theta}$, $J_{\phi\phi}$, $J_{\theta\phi}$, ${\mathbf{J}}_{\bar{\bm{\alpha}}\bar{\bm{\alpha}}}$, and ${\mathbf{j}}_{\phi\bar{\bm{\alpha}}}$ given in (41)–(45), respectively.
@@
Next, we derive matrix ${\mathbf{R}}_{\mathtt{bf}}$ in (36). For the MRT precoder, from (15) and (36), we have
${\mathbf{R}}_{\mathtt{MRT}}=\mathbb{E}\left\{\hat{{\mathbf{H}}}\bm{\Gamma}^{2}\hat{{\mathbf{H}}}^{\scriptscriptstyle\mathsf{H}}+\rho{{\mathbf{v}}}{{\mathbf{v}}}^{\scriptscriptstyle\mathsf{H}}+{{\mathbf{v}}}\bar{\bm{\eta}}^{\scriptscriptstyle\mathsf{T}}\bm{\Gamma}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{H}}}^{\scriptscriptstyle\mathsf{H}}+\hat{{\mathbf{H}}}\bm{\Gamma}\bar{\bm{\eta}}{{\mathbf{v}}}^{\scriptscriptstyle\mathsf{H}}\right\}$
$\quad=\sum_{k=1}^{K}\gamma_{k}\mathbb{E}\left\{\hat{{\mathbf{h}}}_{k}\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}\right\}+\rho{{\mathbf{v}}}{{\mathbf{v}}}^{\scriptscriptstyle\mathsf{H}}=\bm{\xi}_{\mathtt{MRT}}^{\scriptscriptstyle\mathsf{T}}\bm{\gamma}\textbf{{I}}_{N_{\mathtt{t}}}+\rho{{\mathbf{v}}}{{\mathbf{v}}}^{\scriptscriptstyle\mathsf{H}}.$ (97)
Similarly, for the ZF precoder, we have
${\mathbf{R}}_{\mathtt{ZF}}$ $=\underbrace{\mathbb{E}\left\{\hat{{\mathbf{H}}}(\hat{{\mathbf{H}}}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{H}}})^{-1}\bm{\Gamma}^{2}(\hat{{\mathbf{H}}}(\hat{{\mathbf{H}}}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{H}}})^{-1})^{\scriptscriptstyle\mathsf{H}}\right\}}_{\triangleq{\mathbf{A}}}+\rho{{\mathbf{v}}}{{\mathbf{v}}}^{\scriptscriptstyle\mathsf{H}}.$
@@
Let ${\mathbf{A}}={\mathbf{V}}\bm{\Lambda}{\mathbf{V}}^{\scriptscriptstyle\mathsf{H}}$ be the eigenvalue decomposition of ${\mathbf{A}}$. Then ${\mathbf{U}}{\mathbf{V}}\bm{\Lambda}{\mathbf{V}}^{\scriptscriptstyle\mathsf{H}}{\mathbf{U}}^{\scriptscriptstyle\mathsf{H}}={\mathbf{A}}$, implying that any set of mutually orthogonal vectors constitutes valid eigenvectors for ${\mathbf{A}}$. Thus, ${\mathbf{A}}$ must be a scaled identity, i.e., ${\mathbf{A}}=c\textbf{{I}}_{N_{\mathtt{t}}}$, where $c=\frac{1}{N_{\mathtt{t}}(N_{\mathtt{t}}-K)}\sum_{k=1}^{K}\frac{\gamma_{k}}{\xi_{k}}=\bm{\xi}_{\mathtt{ZF}}^{\scriptscriptstyle\mathsf{T}}\bm{\gamma}$, with $\bm{\xi}_{\mathtt{ZF}}$ defined in Lemma 1. As a result, we can write
${\mathbf{R}}_{\mathtt{ZF}}=\bm{\xi}_{\mathtt{ZF}}^{\scriptscriptstyle\mathsf{T}}\bm{\gamma}\textbf{{I}}_{N_{\mathtt{t}}}+\rho{{\mathbf{v}}}{{\mathbf{v}}}^{\scriptscriptstyle\mathsf{H}}.$ (99)
It is observed from (97) and (99) that ${\mathbf{R}}_{\mathtt{ZF}}$ and ${\mathbf{R}}_{\mathtt{MRT}}$ have a similar structure: ${\mathbf{R}}_{\mathtt{bf}}=\bm{\xi}_{\mathtt{bf}}^{\scriptscriptstyle\mathsf{T}}\bm{\gamma}\textbf{{I}}_{N_{\mathtt{t}}}+\rho{{\mathbf{v}}}{{\mathbf{v}}}^{\scriptscriptstyle\mathsf{H}}$, with $\mathtt{bf}\in\{\mathtt{MRT},\mathtt{ZF}\}$. Therefore, $J_{\theta\theta},J_{\phi\phi},J_{\theta\phi},{\mathbf{J}}_{\bar{\bm{\alpha}}\bar{\bm{\alpha}}},$ and $\tilde{J}_{\phi\bar{\bm{\alpha}}}$ for both precoders can be derived in a similar manner and have similar forms. Specifically, from (29), (35), (94)–(97), (99), and noting that $\left\lVert{\mathbf{a}}\right\rVert^{2}=N_{\mathtt{t}}$ and $\left\lVert{\mathbf{b}}\right\rVert^{2}=N_{\mathtt{r}}$, we have
@@
$J_{\theta\theta}$ $=\kappa\left|\alpha\right|^{2}\mathtt{trace}\left\{\left(\dot{{\mathbf{b}}}_{\theta}{\mathbf{a}}^{\scriptscriptstyle\mathsf{H}}+{\mathbf{b}}\dot{{\mathbf{a}}}_{\theta}^{\scriptscriptstyle\mathsf{H}}\right)\left(\bm{\xi}_{\mathtt{bf}}^{\scriptscriptstyle\mathsf{T}}\bm{\gamma}\textbf{{I}}_{N_{\mathtt{t}}}+\rho{{\mathbf{v}}}{{\mathbf{v}}}^{\scriptscriptstyle\mathsf{H}}\right)\right.$
×$\hskip 128.0374pt\times\left.\left({\mathbf{a}}\dot{{\mathbf{b}}}_{\theta}^{\scriptscriptstyle\mathsf{H}}+\dot{{\mathbf{a}}}_{\theta}{\mathbf{b}}^{\scriptscriptstyle\mathsf{H}}\right)\right\}$
@@
$=\kappa\left|\alpha\right|^{2}\Big{(}\bm{\xi}_{\mathtt{bf}}^{\scriptscriptstyle\mathsf{T}}\bm{\gamma}\left(N_{\mathtt{r}}\left\lVert\dot{{\mathbf{a}}}_{\theta}\right\rVert^{2}+N_{\mathtt{t}}\lVert\dot{{\mathbf{b}}}_{\theta}\rVert^{2}\right)$
@@
$\hskip 42.67912pt+\rho{\left({\left\lVert{\mathbf{v}}^{\scriptscriptstyle\mathsf{H}}{\mathbf{a}}\right\rVert^{2}}\lVert\dot{{\mathbf{b}}}_{\theta}\rVert^{2}+N_{\mathtt{r}}{\left\lVert{\mathbf{v}}^{\scriptscriptstyle\mathsf{H}}\dot{{\mathbf{a}}}_{\theta}\right\rVert^{2}}\right)}\Big{)}.$ (100)